OpenStudy (anonymous):

Solve for k. 8k + 2m = 3m + k

OpenStudy (youngstudier):

I'm not sure, sorry.

OpenStudy (youngstudier):

@texaschic101

OpenStudy (texaschic101):

8k + 2m = 3m + k --- subtract k from both sides 8k - k + 2m = 3m --- subtract 2m from both sides 8k - k = 3m - 2m --- combine terms 7k = 1m --- divide by 7 on both sides (m is the same as 1m) k = 1/7m

OpenStudy (olyalhic):

8k+2m=3m+k Add -k to both sides. 8k+2m+−k=k+3m+−k 7k+2m=3m Add -2m to both sides. 7k+2m+−2m=3m+−2m 7k=m Divide both sides by 7. \[\frac{ 7k }{ 7 }=\frac{ m }{ 7 }\] Answer \[k=\frac{ 1 }{ 7 }m\]

OpenStudy (texaschic101):

or you could just put : k = m/7....its the same as k = 1/7m